# Exploring the generalizability of the optimal 0.234 acceptance rate in random-walk Metropolis and parallel tempering algorithms

### v3

@@ 6 Discussion
During the preparation of this revision, the authors used generative AI models (Gemini 2.5 Pro Preview 05-06, Claude 4.0 Sonnet) to assist with coding and manuscript revision. Generative AI was utilized to help generate, debug, and optimize code for the simulations. The new generated code based on PyTorch allows for the use of Graphics Processing Units (GPUs) when running RWM and PT algorithms to significantly speed up simulations and run more simulations at once. For manuscript revision, Generative AI was used to improve phrasing and clarity on text written by the authors. The authors thoroughly reviewed and edited all AI-generated output and take full responsibility for the scientific integrity and final content of this publication.

### v4

@@ Acknowledgements
During the preparation of revisions, the authors used generative AI large language models (Gemini 2.5 Pro Preview 05-06, Claude 4.0 Sonnet) to assist with coding and manuscript revision. Generative AI was utilized to help generate, debug, and optimize code for the simulations. The new generated code based on PyTorch allows for the use of Graphics Processing Units (GPUs) when running RWM and PT algorithms to significantly speed up simulations and run more simulations at once. For manuscript revision, Generative AI was used to improve phrasing and clarity on text written by the authors. The authors thoroughly reviewed and edited all AI-generated output and take full responsibility for the scientific integrity and final content of this publication.